Find an equation of the line containing the given pair of points (-2, -9) and (-6, -7). Please show your steps. Thank you
Answer by checkley71(8403)

FIRST WE FIND THE SLOPE (m)
M=(Y2-Y1)/(X2-X1)=(-7+9)/(-6+2)=2/-4=-1/2 IS THE SLOPE
NOW REPLACE THE X & Y TERMS IN THE LINE EQUATION[Y=mX+b] & SOLVE FOR b.
-9=-1/2*-2+b
-9=1+b
b=-9-1
b=-10 THE Y INTERCEPT.
THUS THE LINE EQUATION IS:
Y=-X/2-10